# Federated Sketching LoRA: On-Device Collaborative Fine-Tuning of Large Language Models

### v2

@@ Abstract
Fine-tuning large language models (LLMs) on devices remains a challenging problem. Recent works have fused low-rank adaptation (LoRA) techniques with federated fine-tuning to mitigate challenges associated with device model sizes and data scarcity. Still, the heterogeneity of resources remains a critical bottleneck: while higher-rank modules generally enhance performance, varying device capabilities constrain LoRA’s feasible rank range. Existing approaches attempting to resolve this issue either lack analytical justification or impose additional computational overhead, leaving a wide gap for efficient and theoretically-grounded solutions. To address these challenges, we propose federated sketching LoRA (FSLoRA), which leverages a sketching mechanism to enable devices to selectively update submatrices of global LoRA modules maintained by the server. By adjusting the sketching ratios, which determine the ranks of the submatrices on the devices, FSLoRA flexibly adapts to device-specific communication and computational constraints. We provide a rigorous convergence analysis of FSLoRA that characterizes how the sketching ratios affect the convergence rate. Through comprehensive experiments on multiple datasets and LLM models, we demonstrate FSLoRA’s performance improvements compared to various baselines. The code is available at https://github.com/wenzhifang/Federated-Sketching-LoRA-Implementation.
@@ 1 Introduction
On-device large language models (LLMs) have recently gained significant attention as a promising complement to cloud-based LLMs [10]. They align with the typical paradigm of LLMs: starting from a base model pre-trained on large-scale datasets to learn general linguistic patterns, semantics, and context, and then undergoing fine-tuning on task-specific data to enhance performance on specialized or domain-specific applications. However, an LLM fine-tuned on a single client device often achieves unsatisfactory performance due to the limited data available on each device. Federated learning [29, 4] has been investigated as a potential solution here, enabling the model to be fine-tuned across a distributed group of clients within the same task domain, without any data sharing.
However, federated learning imposes significant computational and memory costs, as each device must fine-tune the LLM using its local dataset and send updates to the server for model aggregation. Recently, many parameter-efficient fine-tuning methods have been proposed [23, 24, 15] to reduce the cost associated with model adaptation. Among them, low-rank adaptation (LoRA) [15] stands out as a particularly effective approach due to its flexibility. LoRA enables efficient fine-tuning by approximating weight updates $\Delta{\bf W}$ through a low-rank decomposition $\Delta{\bf W}={\bf BA}$, where matrices ${\bf B}$ and ${\bf A}$ contain significantly fewer trainable parameters than the original weight matrix. To support distributed on-device LLM, Zhang et al. [43], Ye et al. [39] incorporated LoRA into conventional federated averaging (FedAvg) [29], significantly reducing the fine-tuning cost by cutting down the number of parameters that need to be synchronized across distributed devices.
Challenges. While integrating federated learning with LoRA reduces the number of trainable parameters via matrix decomposition, computation and communication costs are still forced to increase with the decomposition rank. This poses challenges when complex tasks demand higher-rank LoRA modules, particularly on resource-constrained mobile devices. Furthermore, the heterogeneity in resource availability across distributed devices makes a uniform rank inefficient: a fixed rank $r$ may be too large for some constrained devices, while being too small for more powerful ones, resulting in underutilized resources. Consequently, an approach that reduces computation and communication overhead while adapting LoRA ranks to heterogeneous device capabilities is highly desirable for collaborative fine-tuning of LLMs. Although some existing approaches have attempted to provide a solution here [5, 1, 37], they either lack theoretical justification or impose additional computational overhead, leaving a large gap for an efficient and theoretically-grounded solution. As we discuss in Section 2.2, a comprehensive approach that preserves the analytical and practical benefits of LoRA while enabling heterogeneous on-device fine-tuning under tight resource constraints remains elusive.
@@ 1.1 Contributions
Figure 1: An illustration of our proposed methodology where the server maintains a pair of global LoRA modules while the devices adaptively update submatrices of the global LoRA modules through sketching during each round.
Motivated by these limitations, this work develops a methodology for collaborative on-device LLM fine-tuning that (i) retains the flexibility of LoRA, (ii) provides theoretical convergence guarantees, and (iii) addresses the challenges posed by system heterogeneity and resource constraints across distributed devices. As depicted in Figure 1, our key idea is to introduce a sketching-based LoRA update to the local fine-tuning stage, which allows devices to selectively update a subset of columns and rows of the LoRA modules during each round, reducing the computation and communication consumption. Additionally, our method customizes the fine-tuning process by adjusting the sparsity level of the sketching matrix, i.e., the size of the updated submatrices for each device in each iteration. As we will see, the impact of the introduced sketching mechanism on the overall optimization landscape requires careful modeling consideration, posing additional challenges for the theoretical analysis that we address in this work.
@@
  • We propose federated sketching LoRA (FSLoRA), which leverages a sketching mechanism to enable devices to selectively update submatrices of global LoRA modules maintained by the server. By adjusting the sketching ratios, which determine the ranks of the submatrices on devices, FSLoRA effectively adapts to device-specific communication and computational constraints.
  • We present a rigorous convergence analysis of FSLoRA under non-uniform submatrix update scenarios (i.e., heterogeneous LoRA configurations) across devices, revealing how the sketching ratios affect the convergence rate via scaled smoothness constants. Further, our results show that while increasing the sketching ratios improves convergence theoretically, it also raises communication and computation costs, suggesting a potential trade-off in selecting the sketching ratios.
  • We conduct extensive experiments across multiple datasets and LLM models with diverse parameter settings, demonstrating FSLoRA’s superior performance compared to various baselines in accuracy, training time, and resource utilization. Our ablation studies further validate the effectiveness of the sketching mechanism and the ability of devices to exploit larger global ranks under FSLoRA.
@@ 1.2 Related Works
LoRA-based parameter-efficient fine-tuning: LoRA was introduced in [15] as a parameter-efficient alternative to full model fine-tuning via low-rank matrix approximations. Subsequently, Kalajdzievski [17] proposed rank-stabilized LoRA (rsLoRA), an approach that enhances LoRA’s performance in high-rank scenarios by modifying the scaling factor applied to each low-rank product. Shuttleworth et al. [33] demonstrated that with this design, rsLoRA can approach the performance of full model fine-tuning as the rank of LoRA modules increases. Han et al. [14] introduced a sparse matrix in parallel with LoRA modules to improve the overall adaptation capability. In [25, 28, 38], the authors proposed sequential low-rank adaptation schemes, including ReLoRA and CoLA, to enable high-rank fine-tuning through iterative low-rank updates. However, the works mentioned above focus on centralized scenarios, assuming that the data required for fine-tuning is available at the server.
@@
Collaborative fine-tuning via federated LoRA: Federated LoRA has recently gained attention as a promising approach for efficient and collaborative fine-tuning of LLMs across distributed devices [4, 13]. Sun et al. [34] examined the performance of federated LoRA incorporating differential privacy. To reduce client-server communication overhead, Kuo et al. [22] proposed integrating communication compression with federated LoRA. Meanwhile, Bai et al. [1], Cho et al. [5], Byun and Lee [3], Wang et al. [37], Koo et al. [21] explored the challenges of resource heterogeneity among distributed devices and introduced heterogeneous LoRA as a solution. However, the approaches proposed in [5, 21, 3, 1] lack a theoretical foundation. Moreover, the FlexLoRA method introduced in [1] incurs additional computational overhead due to its reliance on singular value decomposition (SVD). Furthermore, the stack LoRA method proposed in [37] requires the devices to integrate the LoRA modules into the base model, thereby compromising the inherent flexibility of LoRA. Overall, there is still a lack of a systematic and theoretically grounded solution that can effectively tackle the challenges of heterogeneity in collaborative on-device LLM fine-tuning.
@@
Enhancing adaptability via higher-rank LoRA modules: In [15], the authors showed that small ranks can be sufficient for certain tasks; however, they also acknowledge that small rank LoRA modules may not work universally, especially when the downstream task differs significantly from pretraining. Following this, several works explored the effect of increasing the rank in LoRA modules. In a centralized setup, Kalajdzievski [17] and Shuttleworth et al. [33] showed that higher-rank LoRA models can closely approximate full fine-tuning under rsLoRA. In a distributed on-device LLM fine-tuning regime, Bai et al. [1] demonstrated improved performance with larger ranks under FlexLoRA. Similarly, Cho et al. [5] reported that, with proper overfitting control, HeteroLoRA can also benefit from larger ranks. Overall, while small ranks may suffice for simpler tasks or strong base models, higher-rank modules are necessary to compensate for limited backbone capability, such as in on-device LLMs, and to enable effective adaptation to more complex downstream tasks.
@@ 2.1 LoRA-based Federated LLM Fine-tuning
${\bf W}_{0}$ denotes the frozen base model, ${\bf B}\in\mathbb{R}^{m\times r},{\bf A}\in\mathbb{R}^{r\times n}$ are LoRA modules, $N$ denotes the number of devices, $\xi$ denotes a data sample, and $\mathcal{D}_{i}$ is the local dataset on device $i$. $\ell$, $f_{i}$, and $f$ are the sample loss function, the local loss for device $i$, and the global loss, respectively.
Problem (1) aligns with the conventional federated optimization formulation, which thus can be solved using the FedAvg algorithm. Based on the FedAvg framework, Zhang et al. [43] developed federated LoRA, which applies a uniform rank $r$ across all devices, overlooking system heterogeneity. This one-size-fits-all approach leads to resource mismatches, where computationally constrained devices may struggle, while more powerful devices remain underutilized with a fixed rank.
@@ 2.2 Aren’t the Existing Solutions Good Enough?
To address this issue, researchers have proposed heterogeneous federated LoRA approaches, where devices maintain non-uniform LoRA modules with varying ranks. They also introduce mechanisms to overcome the challenges of directly aggregating matrices with different dimensions. However, these methods often lack theoretical foundation or incur additional computational and memory overhead.
HeteroLoRA [5] lets the server pad the updates from the devices with smaller ranks to match the size of the largest rank during aggregation. During model dissemination, devices receive a truncated version of the global LoRA modules from the server. Although easy to implement, HeteroLoRA is primarily heuristic in nature and lacks a rigorous theoretical foundation, potentially limiting its performance, as we will see in Section 5.
FlexLoRA [1] requires the server to collect the individual LoRA matrices $\mathbf{B}_{i}$ and $\mathbf{A}_{i}$ from the devices and then computes their product $\mathbf{B}_{i}\mathbf{A}_{i}$. To support the initialization of non-uniform LoRA modules, the server applies truncated SVD to the averaged product $\frac{1}{N}\sum_{i=1}^{N}\mathbf{B}_{i}\mathbf{A}_{i}$. However, this approach introduces extra computational and memory overhead on the server due to truncated SVD, and the associated error can limit the performance as demonstrated in Section 5.
FedStackLoRA [37] introduces a stacking mechanism where the server concatenates LoRA modules from the devices. The concatenated matrices are then sent back to the devices, which compute their product and merge it into the base model before initializing new LoRA modules for the next fine-tuning round. However, this approach increases communication complexity linearly with the number of devices, imposes higher computation and memory demands on the devices, and compromises LoRA’s flexibility to support multiple adapters for different tasks.
More detailed comparisons on computation, memory, and communication are presented in Appendix A. In summary, a theoretically-grounded solution that preserves LoRA’s benefits while effectively addressing system heterogeneity and the constraints of resource-limited devices remains lacking.
@@ 3 Federated Sketching LoRA
Motivated by the limitations of existing methods, we propose a new federated LoRA reformulation. Building on this foundation, we develop FSLoRA, a resource-adaptive algorithm that preserves LoRA’s flexibility while accommodating for heterogeneous device capabilities.
@@ 3.1 Our Formulation
where ${\bf B}\in\mathbb{R}^{m\times r},{\bf A}\in\mathbb{R}^{r\times n}$ are LoRA modules, $f_{i}^{\mathcal{S}}$ is the local loss function at device $i$ with sketching, and ${\bf S}$ denotes a sketching matrix randomly sampled from the diagonal matrix set $\mathcal{S}_{i}=\mathcal{S}(r,k_{i})$. The set $\mathcal{S}(r,k_{i})$ comprises diagonal matrices of size $r\times r$ with exactly $k_{i}$ non-zero entries. The formal definition of $\mathcal{S}(r,k)$ is provided below:
@@
where $\mathbf{e}_{1},\dots,\mathbf{e}_{r}\in\mathbb{R}^{r}$ are standard unit basis vectors and index set $\mathcal{I}$ is a random subset of $[r]\coloneqq\{1,2,\ldots,r\}$ sampled uniformly from all subsets of $[r]$ with cardinality $k$.
With $\mathbf{S}$ being a matrix sampled from $\mathcal{S}_{i}$, we have $\mathbf{BSA}=\frac{r}{k_{i}}\sum_{j\in\mathcal{I}_{i}}\mathbf{B}\mathbf{e}_{j}\mathbf{e}_{j}^{\top}\mathbf{A},$ where $\mathcal{I}_{i}$ corresponds to the index set of non-zero entries of $\mathbf{S}$. $\mathbf{B}\mathbf{e}_{j}$ extracts the $j$-th column of $\mathbf{B}$ while $\mathbf{e}_{j}^{\top}\mathbf{A}$ extracts the $j$-th row of $\mathbf{A}$. In other words, only $k_{i}$ columns and rows in the LoRA modules $\bf B$ and $\bf A$ are activated by the sketching matrix in the loss $\ell({\bf W}_{0}+{\bf B}\bf S{\bf A},\xi)$ at device $i$. On the other hand, the sketching matrix $\bf S$ satisfies $\mathbb{E}_{\mathbf{S}\sim\mathcal{S}_{i}}[\mathbf{S}]=\mathbf{I}_{r}$ where $\mathbf{I}_{r}$ is a $r$-dimensional identity matrix. Based upon this property, ${\bf W}_{0}+{\bf B}{\bf S}{\bf A}$ can be treated as an unbiased estimate of ${\bf W}_{0}+\bf B{\bf A}$.
Intuition: A larger rank allows LoRA modules to be more expressive, leading to better performance [1, 17, 33]. However, resource-constrained devices cannot afford the computational or communication demands of large-rank modules. Our formulation (2) leverages the sketching matrix to balance the expressiveness of high-rank LoRA modules with the resource constraints of different devices. With the sketching mechanism introduced, the local gradients with respect to the LoRA modules on the devices will exhibit structured sparsity. By adjusting the sketching ratio $k_{i}/r$, we can tailor the sparsity of the gradient to match the capabilities of each device, ensuring affordable training while maintaining performance across heterogeneous systems, as elaborated in the following subsection. Overall, compared to the original objective in (1), our formulation offers a more resource-adaptive and flexible framework, tailored to address the diverse capabilities of heterogeneous devices.

### v3

# Federated Sketching LoRA: A Flexible Framework for Heterogeneous Collaborative Fine-Tuning of LLMs

@@ Abstract
Fine-tuning large language models (LLMs) on resource-constrained clients remains a challenging problem. Recent works have fused low-rank adaptation (LoRA) techniques with federated fine-tuning to mitigate challenges associated with client model sizes and data scarcity. Still, the heterogeneity of resources remains a critical bottleneck: while higher-rank modules generally enhance performance, varying client capabilities constrain LoRA’s feasible rank range. Existing approaches attempting to resolve this issue either lack analytical justification or impose additional computational overhead, leaving a wide gap for efficient and theoretically-grounded solutions. To address these challenges, we propose federated sketching LoRA (FSLoRA), which leverages a sketching mechanism to enable clients to selectively update submatrices of global LoRA modules maintained by the server. By adjusting the sketching ratios, which determine the ranks of the submatrices on the clients, FSLoRA flexibly adapts to client-specific communication and computational constraints. We provide a rigorous convergence analysis of FSLoRA that characterizes how the sketching ratios affect the convergence rate. Through comprehensive experiments on multiple datasets and LLM models, we demonstrate FSLoRA’s performance improvements compared to various baselines. The code is available at https://github.com/wenzhifang/Federated-Sketching-LoRA-Implementation.
@@ 1 Introduction
Lightweight client-side large language models (LLMs) have recently gained significant attention as a promising complement to cloud-based LLMs [13]. They align with the typical paradigm of LLMs: starting from a base model pre-trained on large-scale datasets to learn general linguistic patterns, semantics, and context, and then undergoing fine-tuning on task-specific data to enhance performance on specialized or domain-specific applications. However, an LLM fine-tuned on a single client often achieves unsatisfactory performance due to the limited data. Federated learning [30, 7] has been investigated as a potential solution here, enabling the model to be fine-tuned across a group of distributed clients within the same task domain, without any raw data sharing.
However, federated LLM fine-tuning is costly in both computation and communication due to the massive parameter volume. Importantly, many parameter-efficient fine-tuning methods have been proposed [26, 27, 18] to reduce the model adaptation cost. Among them, low-rank adaptation (LoRA) [18] stands out as a particularly effective approach due to its flexibility. In particular, LoRA enables efficient fine-tuning by approximating weight updates $\Delta{\bf W}$ through a low-rank decomposition $\Delta{\bf W}={\bf BA}$, where matrices $\bf B$} and $\bf A$} contain significantly fewer trainable parameters than the original weight matrix. Building on this foundation, recent works have combined LoRA with federated averaging (FedAvg) [47, 43], showing that federated LoRA significantly reduce the training overhead.
Challenges. While incorporating LoRA into federated LLM fine-tuning reduces the number of trainable parameters, computation and communication costs are still forced to increase with the LoRA rank. This poses challenges when complex tasks demand higher-rank LoRA modules, particularly on resource-constrained clients. Furthermore, the heterogeneity in resource availability across distributed clients makes a uniform rank adopted in federated LoRA inefficient: a fixed rank $r$ may be too large for some constrained clients, while being too small for more powerful ones, resulting in underutilized resources. Consequently, an approach that further reduces computation and communication overhead while adapting LoRA ranks to heterogeneous client capabilities is highly desirable. Although some existing approaches have attempted to provide a solution [8, 1, 42], they either lack theoretical justification or impose additional computational overhead, leaving a gap for an efficient and theoretically-grounded solution. As we discuss in Section 2.2, a comprehensive approach that preserves the analytical and practical benefits of LoRA while enabling heterogeneous collaborative fine-tuning under tight resource constraints remains elusive.
@@ 1.1 Contributions
Figure 1: An illustration of our proposed methodology where the server maintains a pair of global LoRA modules while the clients adaptively update submatrices of the global LoRA modules through sketching during each round.
Motivated by these limitations, this work develops a methodology for efficient federated LLM fine-tuning that (i) retains the flexibility of LoRA, (ii) provides theoretical convergence guarantees, and (iii) addresses the challenges posed by heterogeneous and constrained resources across distributed clients. As depicted in Figure 1, our key idea is to introduce a sketching-based LoRA update to the local fine-tuning, which allows clients to selectively update a subset of columns and rows of the LoRA modules during each round, reducing the computation and communication consumption. Additionally, our method customizes the fine-tuning process by adjusting the sparsity level of the sketching matrix, i.e., the size of the updated submatrices for each client in each iteration. As we will see, the impact of the introduced sketching mechanism on the overall optimization landscape requires careful modeling consideration, posing additional challenges for the theoretical analysis that we address in this work.
@@
  • We propose federated sketching LoRA (FSLoRA), which leverages a sketching mechanism to enable clients to selectively update submatrices of global LoRA modules maintained by the server. By adjusting the sketching ratios, which determine the ranks of the submatrices on clients, FSLoRA effectively adapts to client-specific communication and computational constraints.
  • We present a rigorous convergence analysis of FSLoRA under non-uniform submatrix update scenarios (i.e., heterogeneous LoRA configurations) across clients, revealing how the sketching ratios affect the convergence rate via scaled smoothness constants. Further, our results show that while increasing the sketching ratios improves convergence theoretically, it also raises communication and computation costs, suggesting a potential trade-off in selecting the sketching ratios.
  • We conduct extensive experiments across multiple datasets and LLM models with diverse parameter settings, demonstrating FSLoRA’s superior performance compared to various baselines in accuracy, training time, and resource utilization. Our ablation studies further validate the effectiveness of the sketching mechanism and the ability of clients to exploit larger global ranks under FSLoRA.
@@ 1.2 Related Works
Collaborative fine-tuning via federated LoRA: Federated LoRA is an efficient approach for collaborative LLM fine-tuning among distributed clients [7, 38, 17]. Building on this foundation, Kuo et al. [25] proposed integrating communication compression with federated LoRA to further reduce communication overhead. Meanwhile, Bai et al. [1], Cho et al. [8], Byun and Lee [5], Wang et al. [42], Koo et al. [24] explored the challenges of resource heterogeneity across distributed clients and introduced heterogeneous LoRA as a solution. However, the approaches proposed in [8, 24, 5, 1] lack a theoretical foundation. Moreover, the FlexLoRA method introduced in [1] incurs additional computational overhead due to its reliance on singular value decomposition (SVD). Furthermore, the FLoRA algorithm proposed in [42] requires the clients to merge the LoRA modules into the base model, thereby compromising the inherent flexibility of LoRA. Overall, there is still a lack of a systematic and theoretically grounded solution that can effectively tackle heterogeneous collaborative LLM fine-tuning.
@@
Enhancing adaptability via higher-rank LoRA modules: The foundational study by Hu et al. [18] demonstrated that small ranks can be sufficient for certain tasks; however, they also acknowledge that small rank LoRA modules may not work universally, especially when the downstream task differs significantly from pretraining. Following this, several works explored the effect of increasing the rank in LoRA modules. In a centralized setup, Kalajdzievski [20] and Shuttleworth et al. [37] showed that higher-rank LoRA models can closely approximate full fine-tuning under rsLoRA. In a federated LLM fine-tuning regime, Bai et al. [1] demonstrated improved performance with larger ranks under FlexLoRA. Similarly, Cho et al. [8] reported that, with proper overfitting control, HeteroLoRA can also benefit from larger ranks. Overall, while small ranks may suffice for simpler tasks or strong base models, higher-rank modules are necessary to compensate for limited backbone capability, such as in lightweight LLMs, and to enable effective adaptation to more complex downstream tasks.
Sketching-based optimization: Sketching is an efficient technique for mitigating the complexity of high-dimensional optimization, with its earliest applications in least-squares regression [35, 41]. Beyond this, gradient sketching has been employed to construct preconditioners for gradient descent methods [16]. Building on these foundations, recent work has applied sketching to distributed optimization. In particular, Charalambides and Mazumdar [6] proposed hybrid local-global sketching for distributed least-squares, while Demidovich et al. [11] developed a distributed sparsified training framework based on sketching. Shrivastava et al. [36] demonstrated that sketching substantially reduces communication in distributed training of overparameterized deep models without sacrificing accuracy. More recently, Nicolas et al. [32] investigated sketching-based differential privacy and demonstrated its compatibility with secure aggregation. Despite these advances, sketching strategies tailored to structured low-rank adaptation modules such as LoRA remain largely unexplored.
@@ 2.1 LoRA-based Federated LLM Fine-tuning
where ${\bf W}_{0}$ denotes the frozen base model, ${\bf B}\in\mathbb{R}^{m\times r},{\bf A}\in\mathbb{R}^{r\times n}$ are LoRA modules, $N$ denotes the number of clients, $\xi$ denotes a data sample, and $\mathcal{D}_{i}$ is the local dataset on client $i$. $\ell$, $f_{i}$, and $f$ are the sample loss function, the local loss for client $i$, and the global loss, respectively.
Problem (1) aligns with the conventional federated optimization formulation, which thus can be solved using the FedAvg algorithm. Based on the FedAvg framework, Zhang et al. [47] developed federated LoRA, which applies a uniform rank $r$ across all clients, overlooking resource heterogeneity. This one-size-fits-all approach leads to resource mismatches, where computationally constrained clients may struggle, while more powerful clients remain underutilized with a fixed rank.
@@ 2.2 Aren’t the Existing Solutions Good Enough?
To address this issue, researchers have proposed heterogeneous federated LoRA approaches, where clients maintain non-uniform LoRA modules with varying ranks. They also introduce mechanisms to overcome the challenges of directly aggregating matrices with different dimensions. However, these methods often lack theoretical foundation or incur additional computational and memory overhead.
HeteroLoRA [8] lets the server pad the updates from the clients with smaller ranks to match the size of the largest rank during aggregation. During model dissemination, clients receive a truncated version of the global LoRA modules from the server. Although easy to implement, HeteroLoRA is primarily heuristic in nature and lacks a rigorous theoretical foundation, potentially limiting its performance, as we will see in Section 5.
FlexLoRA [1] requires the server to collect the individual LoRA matrices $\mathbf{B}_{i}$ and $\mathbf{A}_{i}$ from the clients and then computes their product $\mathbf{B}_{i}\mathbf{A}_{i}$. To support the initialization of non-uniform LoRA modules, the server applies truncated SVD to the averaged product $\frac{1}{N}\sum_{i=1}^{N}\mathbf{B}_{i}\mathbf{A}_{i}$. However, this approach introduces extra computational and memory overhead on the server due to truncated SVD, and the associated error can limit the performance as demonstrated in Section 5.
FLoRA [42] introduces a stacking mechanism where the server concatenates LoRA modules from the clients. The concatenated matrices are then sent back to the clients, which compute their product and merge it into the base model before initializing new LoRA modules for the next fine-tuning round. However, this approach increases communication complexity linearly with the number of clients, imposes higher computation and memory demands on the clients, and compromises LoRA’s flexibility to support multiple adapters for different tasks.
More detailed comparisons on computation, memory, and communication are presented in Appendix A. In summary, a theoretically-grounded solution that preserves LoRA’s benefits while effectively addressing resource heterogeneity across distributed clients remains lacking.
@@ 3 Federated Sketching LoRA
Motivated by the limitations of existing methods, we propose a new federated LoRA reformulation. Building on this foundation, we develop FSLoRA, a heterogeneous LoRA algorithm that preserves LoRA’s flexibility while accommodating client resource heterogeneity.
@@ 3.1 Our Formulation
where ${\bf B}\in\mathbb{R}^{m\times r},{\bf A}\in\mathbb{R}^{r\times n}$ are LoRA modules, $f_{i}^{\mathcal{S}}$ is the local loss function at client $i$ with sketching, and $\bf S$} denotes a sketching matrix randomly sampled from the diagonal matrix set $\mathcal{S}_{i}=\mathcal{S}(r,k_{i})$. The set $\mathcal{S}(r,k_{i})$ comprises diagonal matrices of size $r\times r$ with exactly $k_{i}$ non-zero entries. The formal definition of $\mathcal{S}(r,k)$ is provided below:
@@
where $\mathbf{e}_{1},\dots,\mathbf{e}_{r}\in\mathbb{R}^{r}$ are standard unit basis vectors and index set $\mathcal{I}$ is a random subset of $r]\coloneqq\{1,2,\ldots,r\}$ sampled uniformly from all subsets of $[r]$ with cardinality $k$.
With $\mathbf{S}$ being a matrix sampled from $\mathcal{S}_{i}$, we have $\mathbf{BSA}=\frac{r}{k_{i}}\sum_{j\in\mathcal{I}_{i}}\mathbf{B}\mathbf{e}_{j}\mathbf{e}_{j}^{\top}\mathbf{A},$ where $\mathcal{I}_{i}$ corresponds to the index set of non-zero diagonal entries of $\mathbf{S}$. $\mathbf{B}\mathbf{e}_{j}$ extracts the $j$-th column of $\mathbf{B}$ while $\mathbf{e}_{j}^{\top}\mathbf{A}$ extracts the $j$-th row of $\mathbf{A}$. In other words, only $k_{i}$ columns and rows in the LoRA modules $\bf B$ and $\bf A$ are activated by the sketching matrix in the loss $\ell({\bf W}_{0}+{\bf B}\bf S{\bf A},\xi)$ at client $i$. On the other hand, the sketching matrix $\bf S$ satisfies $\mathbb{E}_{\mathbf{S}\sim\mathcal{S}_{i}}[\mathbf{S}]=\mathbf{I}_{r}$ where $\mathbf{I}_{r}$ is a $r$-dimensional identity matrix. Based upon this property, ${\bf W}_{0}+{\bf B}{\bf S}{\bf A}$ can be treated as an unbiased estimate of ${\bf W}_{0}+\bf B{\bf A}$.
Intuition: A larger rank allows LoRA modules to be more expressive, leading to better performance [1, 20, 37]. However, resource-constrained clients cannot afford the computational or communication demands of large-rank modules. Our formulation (2) leverages the sketching matrix to balance the expressiveness of high-rank LoRA modules with the resource constraints of different clients. With the sketching mechanism introduced, the local gradients with respect to the LoRA modules on the clients will exhibit structured sparsity. By adjusting the sketching ratio $k_{i}/r$, we can tailor the sparsity of the gradient to match the capabilities of each client, ensuring affordable training while maintaining performance across heterogeneous systems, as elaborated in the following subsection. Overall, compared to (1), our formulation offers a more flexible framework, tailored to address the diverse capabilities of heterogeneous clients.